Grade 11Physical Chemistry

how much should the pressure be increased in order to decrease the volume of a gas by 5% at a constant tempreture?

Raviraj singh
11 Years agoGrade 11

2 Answers

k durgapavani
11 Years ago
AT CONSTANT TEMPERATURE PRESSURE AND VOLUME ARE INVERESLY PROPORTIONAL.SO WHEN THE VOLUME OF A GAS IS DECREASED BY 5% PRESSURE SHOULD BE INCREASED BY 5%
Shreya ojha
7 Years ago
P1V1=P2V2
Volume decreased by 5%=100-5/100=95/100
V2=95/100 V1
Putting the values we get:-
P1V1=P2×95/100V1
P2=100/95=1.05
So pressure should be increased by 5.26%
P2=100/95=